michelle is 19 years older than her brother doug. 6 years fro now, michelle will be twice as old as doug. how old is each person now?
Answer by scott8148(6628)

"michelle is 19 years older than her brother doug" ___ m = d + 19

"6 years fro now, michelle will be twice as old as doug" ___ m + 6 = 2(d + 6)

substituting the equivalent value for m ___ (d + 19) + 6 = 2d + 12

13 = d

substituting ___ m = (13) + 19 = 32